# DISCO: Diversifying Sample Condensation for Efficient Model Evaluation

### v1

@@ Appendix A Mutual Information and Jensen-Shannon Divergence
A.0.1 Setup.
Denote a set of source models as $\mathcal{F}=\{f^{1},\ldots,f^{M}\}$. Let $m\sim\mathrm{Unif}\{1,\dots,M\}$ be the index of a uniformly chosen model and $S(m)=S(f^{m},\mathcal{D})$ denote any statistic that is a deterministic function of $m$ computed on $\mathcal{D}$ (e.g. accuracy on $\mathcal{D}$). Let $f_{c}^{m}(x_{i})\in[0,1]$ be the predictive probability for class $c$ of model $f^{m}$ on input $x_{i}$. We write $\widehat{y}_{i}^{m}$ for the categorical random variable following $\operatorname{Cat}(f_{1}^{m}(x_{i}),\ldots,f_{C}^{m}(x_{i}))$ (i.e., given $m$, the prediction on datapoint $i$ is a categorical random variable $\widehat{y}_{i}^{m}$). Define ensemble mean prediction to be $\widebar{f}_{c}(x_{i}):=\mathbb{E}_{m}[f^{m}_{c}(x_{i})]$ for each class $c$ and define corresponding prediction random variable as $\widehat{y}_{i}$ following $\operatorname{Cat}(\widebar{f}_{1}(x_{i}),\ldots,\widebar{f}_{C}(x_{i}))$.
As a consequence, the ensemble mean prediction random variable $\widehat{y}_{i}$ coincides with marginal distribution of $\widehat{y}_{i}^{m}$:
$\Pr(\widehat{y}_{i}=c)\;=\;\mathbb{E}_{m\sim\operatorname{Unif}[M]}\!\left[\Pr(\widehat{y}_{i}=c|m)\right]\;=\;\frac{1}{M}\sum_{m=1}^{M}\Pr(\widehat{y}_{i}^{m}=c).$
Definition 1 (Jensen–Shannon divergence).
Generalised Jensen-Shannon Divergence for multiple distributions (Fuglede & Topsoe, 2004):
$\mathrm{JSD}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)=\frac{1}{M}\sum_{m=1}^{M}D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i})=H(\widehat{y}_{i})-\tfrac{1}{M}\sum_{m=1}^{M}H(\widehat{y}_{i}^{m}).$
A.0.2 Assumptions.
Assumption A1 (Deterministic predictions).
Conditional on $m$, each prediction $\widehat{y}_{i}^{m}$ is fully determined by $m$ (or more generally, any residual randomness is independent across $i$ and independent of $m$).
@@
A.0.3 Proposition.
@@
Under Assumptions A1A2, if $S(m)$ is injective, then
$\mathrm{MI}_{m,\widehat{y}_{i}}\left(S(m);\widehat{y}_{i}\right)\;=\;{H}(\widehat{y}_{i})\;-\tfrac{1}{M}\sum_{m=1}^{M}H(\widehat{y}_{i}^{m})\;=:\;\mathrm{JSD}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right).$
where ${H}(\cdot)$ is entropy, $\mathrm{MI}(\cdot)$ is mutual information.
@@
$\widehat{y}_{i}^{m}\;\longleftrightarrow\;m\;\longleftrightarrow\;S(m).$
@@
$\mathrm{MI}\big(S(m);\widehat{y}_{i}\big)=\mathrm{MI}\big(m;\widehat{y}_{i}\big).$
@@
$\mathrm{MI}(m;\widehat{y}_{i})={H}(\widehat{y}_{i})-\mathbb{E}_{m}\!\left[{H}\left(\widehat{y}_{i}|m\right)\right].$
@@
$\mathbb{E}_{m}\big[{H}(\widehat{y}_{i}|m)\big]=\tfrac{1}{M}\sum_{m=1}^{M}{H}(\widehat{y}_{i}^{m}).$
@@
$\mathrm{MI}(m;\widehat{y}_{i})={H}(\widehat{y}_{i})-\tfrac{1}{M}\sum_{m=1}^{M}{H}(\widehat{y}_{i}^{m})=:\mathrm{JSD}\left(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\right).$
Appendix B Bounds for Jensen-Shannon Divergence (JSD) via Predictive Diversity Score (PDS)
In this section, we show that JSD (Definition 1) is bounded quadratically below and linearly above by PDS. We first relate JSD to total variation (§ B.1), then show total variation is monotone in PDS (§ B.2), and then combine these results in § B.3.
B.1 Bounds for JSD via Total Variation (TV)
We begin by showing that JSD is bounded quadratically below and linearly above by total variation. We first introduce the setup with required definitions (§ B.1.1), then prove the proposition (§ B.1.2).
B.1.1 Setup.
Use notation from Appendix A. Let $\widehat{y}_{i}^{m}$ for $m=1,\ldots,M$ be categorical random variables on $K$ classes, and let $\widehat{y}_{i}$ be the ensemble mean prediction random variable defined as the mixture
$\Pr(\widehat{y}_{i}=c)\;=\;\frac{1}{M}\sum_{m=1}^{M}\Pr(\widehat{y}_{i}^{m}=c).$
@@ B.1 Bounds for JSD via Total Variation (TV)
For categorical random variables $X,Y$ on the same support,
$\mathrm{TV}(X,Y)=\tfrac{1}{2}\sum_{c}|\Pr(X=c)-\Pr(Y=c)|.$
B.1.2 Proposition.
@@
For any $M\geq 2$ categorical random variables $\{\widehat{y}_{i}^{m}\}_{m=1}^{M}$ with ensemble mean $\widehat{y}_{i}$,
$\frac{2}{\ln 2}\cdot\frac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i})^{2}\;\leq\;\mathrm{JSD}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)\;\leq\;\frac{M}{M-1}\,\log M\cdot\frac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}).$
@@
Lower bound. From Definition 1,
$\mathrm{JSD}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)=\frac{1}{M}\sum_{m=1}^{M}D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i}).$
By Pinsker’s inequality (e.g. Equation 1 in (Sason, 2015)),
$D_{\mathrm{KL}}(X\|Y)\;\geq\;\frac{2}{\ln 2}\,\mathrm{TV}(X,Y)^{2}.$
Substituting $X=\widehat{y}_{i}^{m}$, $Y=\widehat{y}_{i}$ and averaging over $m$ yields the lower bound.
Upper bound. Fix $m$. For each class $c$, write
$\Pr(\widehat{y}_{i}=c)=\alpha\Pr(\widehat{y}_{i}^{m}=c)+(1-\alpha)\zeta(c),\quad\alpha=\tfrac{1}{M},\quad\zeta(c)=\tfrac{1}{M-1}\sum_{s\neq m}\Pr(\widehat{y}_{i}^{s}=c).$
Define $t(c)=\zeta(c)/\Pr(\widehat{y}_{i}^{m}=c)$ when $\Pr(\widehat{y}_{i}^{m}=c)>0$ (set $t(c)=+\infty$ if $\Pr(\widehat{y}_{i}^{m}=c)=0,\zeta(c)>0$). Then $\mathbb{E}_{\widehat{y}_{i}^{m}}[t]=1$ and
$D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i})=\mathbb{E}_{\widehat{y}_{i}^{m}}\!\left[-\log\big(\alpha+(1-\alpha)t\big)\right].$
Let $g(u)=-\log(\alpha+(1-\alpha)u)$, $u\geq 0$. Then $g$ is convex, decreasing, with $g(1)=0$, $g(0)=\log(1/\alpha)=\log M$. By convexity,
$g(u)\;\leq\;(1-u)g(0)=(1-u)\log M.$
@@
$D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i})\leq\log M\cdot\mathbb{E}_{\widehat{y}_{i}^{m}}\!\big[(1-t)\big]\leq\log M\cdot\mathbb{E}_{\widehat{y}_{i}^{m}}\!\big[(1-t)_{+}\big].$
@@
$\mathbb{E}_{\widehat{y}_{i}^{m}}[(1-t)_{+}]=\sum_{c}\Pr(\widehat{y}_{i}^{m}=c)\max\{0,1-\zeta(c)/\Pr(\widehat{y}_{i}^{m}=c)\}=\sum_{c}(\Pr(\widehat{y}_{i}^{m}=c)-\zeta(c))_{+}.$
@@
$\sum_{c}(\Pr(\widehat{y}_{i}^{m}=c)-\zeta(c))_{+}=\mathrm{TV}(\widehat{y}_{i}^{m},\zeta).$
Finally, since $\Pr(\widehat{y}_{i}=c)=\alpha\Pr(\widehat{y}_{i}^{m}=c)+(1-\alpha)\zeta(c)$, one has
$\mathrm{TV}(\widehat{y}_{i}^{m},\zeta)=\tfrac{M}{M-1}\,\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}).$
@@
$D_{\mathrm{KL}}(\widehat{y}_{i}^{m}\|\widehat{y}_{i})\;\leq\;\tfrac{M}{M-1}\,\log M\cdot\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}).$
@@
B.2 Bounds for Total Variation via Predictive Diversity Score
We next show that total variation is monotone in PDS. We introduce the setup with definitions and lemmas (§ B.2.1), then prove the proposition (§ B.2.2).
B.2.1 Setup.
Fix a class $c$. Let $X_{m}=\Pr(\widehat{y}_{i}^{m}=c)$ and $\mu=\Pr(\widehat{y}_{i}=c)$. Define:
@@ B.2 Bounds for Total Variation via Predictive Diversity Score
$\operatorname{PDS}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)=\sum^{K}_{c=1}\max_{m}\Pr(\widehat{y}_{i}^{m}=c).$
@@
Use notation from Appendix B.2.1. For each class $c$, if at most $z$ models satisfy $X_{m}>\mu$, then
@@
$E=\sum_{c=1}^{K}E_{c},\qquad U=\sum_{c=1}^{K}U_{c}=\tfrac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}).$
@@
B.3 Final Sandwich Inequality
Finally, we combine results from § B.1.1 and § B.2 to show that JSD is bounded quadratically below and linearly above by PDS.
@@ B.3 Final Sandwich Inequality
$\frac{2}{M^{2}\ln 2}\,(\mathrm{PDS}\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)-1)^{2}\;\leq\;\mathrm{JSD}\!\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)\;\leq\;\frac{M}{M-1}\log M\cdot(\mathrm{PDS}\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)-1).$
@@
$U\;=\;\sum_{c=1}^{K}\frac{1}{2M}\sum_{m=1}^{M}|\Pr(\widehat{y}_{i}^{m}=c)-\Pr(\widehat{y}_{i}=c)|=\frac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(\widehat{y}_{i}^{m},\widehat{y}_{i}),$
and
$E\;=\;\sum_{c=1}^{K}\Big(\max_{m}\Pr(\widehat{y}_{i}^{m}=c)-\Pr(\widehat{y}_{i}=c)\Big)=\mathrm{PDS}\big(\widehat{y}_{i}^{1},\ldots,\widehat{y}_{i}^{M}\big)-1.$
@@
Appendix C Vision results
We introduce the setup and present results.

### v2

@@ Appendix G Mutual Information and Jensen-Shannon Divergence
G.0.1 Setup.
Let $m\sim\mathrm{Unif}\{1,\dots,M\}$ be the index of a uniformly chosen model. Given $m$, the prediction on datapoint $i$ has categorical law $P_{\widehat{Y}_{i}\mid m}$. Define the ensemble mean distribution
$P_{\widehat{Y}_{i}}\;=\;\mathbb{E}_{m\sim\operatorname{Unif}[M]}\!\left[P_{\widehat{Y}_{i}\mid m}\right]\;=\;\frac{1}{M}\sum_{m=1}^{M}P_{\widehat{Y}_{i}\mid m}.$
Let $S(m)$ denote any statistic that is a deterministic function of $m$ computed on $\mathcal{D}$ (e.g. accuracy on $\mathcal{D}$).
G.0.2 Assumptions.
@@
We note that Assumption A1 does not assume a uniform prior over the source models $f^{1},\ldots,f^{M}$. It only assumes that the model index is drawn uniformly, $m\sim\mathrm{Unif}{1,\ldots,M}$. If the prior over models is non-uniform, we can replicate models proportionally to their sampling probabilities. In this case, the index distribution can be made uniform without changing the resulting model sampling outcomes.
Assumption A2 (Deterministic predictions).
Conditional on $m$, each prediction $\widehat{Y}_{i}$ is fully determined by $m$ (or more generally, any residual randomness is independent across $i$ and independent of $m$).
G.0.3 Proposition.
@@
Under Assumptions A2A1, if $S(m)$ is injective, then
$\mathrm{MI}_{m,\widehat{Y}}\!\left(S(m)\,;\,\widehat{Y}_{i}\right)\;=\;\mathcal{H}_{\widehat{Y}_{i}}\!\left(P_{\widehat{Y}_{i}}\right)\;-\mathbb{E}_{m\sim\operatorname{Unif}[M]}\!\left[\mathcal{H}_{\widehat{Y}_{i}}\!\left(P_{\widehat{Y}_{i}\mid m}\right)\right]\;=:\;\mathrm{JSD}\big(\{P_{\widehat{Y}_{i}\mid m}\}_{m=1}^{M}\big).$
@@
$\widehat{Y}_{i}\;\longleftrightarrow\;m\;\longleftrightarrow\;S(m).$
@@
$I\big(S(m);\widehat{Y}_{i}\big)=I\big(m;\widehat{Y}_{i}\big).$
@@
$I(m;\widehat{Y}_{i})=\mathcal{H}_{\widehat{Y}_{i}}\!\left(P_{\widehat{Y}_{i}}\right)-\mathbb{E}_{m}\!\left[\mathcal{H}_{\widehat{Y}_{i}}\!\left(P_{\widehat{Y}_{i}\mid m}\right)\right].$
Marginal distribution (using Assumption A1):
$P_{\widehat{Y}_{i}}=\sum_{m=1}^{M}\Pr(m)\,P_{\widehat{Y}_{i}\mid m}=\tfrac{1}{M}\sum_{m=1}^{M}P_{\widehat{Y}_{i}\mid m}.$
Thus $\mathcal{H}_{\widehat{Y}_{i}}(P_{\widehat{Y}_{i}})$ is the entropy of the mixture distribution.
@@
$\mathbb{E}_{m}\big[\mathcal{H}_{\widehat{Y}_{i}}(P_{\widehat{Y}_{i}\mid m})\big]=\tfrac{1}{M}\sum_{m=1}^{M}\mathcal{H}_{\widehat{Y}_{i}}(P_{\widehat{Y}_{i}\mid m}).$
@@
$I(m;\widehat{Y}_{i})=\mathcal{H}_{\widehat{Y}_{i}}(P_{\widehat{Y}_{i}})-\tfrac{1}{M}\sum_{m=1}^{M}\mathcal{H}_{\widehat{Y}_{i}}(P_{\widehat{Y}_{i}\mid m})=:\mathrm{JSD}\big(\{P_{\widehat{Y}_{i}\mid m}\}_{m=1}^{M}\big).$
We note that JSD and, as a consequence, DISCO predictor directly depend on the heterogeneity of source models. This heterogeneity is captured by how distinguishable the model-conditional predictive distributions $P_{\widehat{Y}_{i}\mid m}$ are from their mixture, as measured by their average KL divergence to the ensemble mean. The larger this KL divergence, the higher the JSD. According to Proposition 3, a larger JSD is equivalent to higher mutual information between outputs and benchmark accuracy, which leads to better performance of the DISCO predictor. Conversely, if this KL divergence is small, the JSD is also small. In particular, if we have many copies of the same model, then JSD as well as mutual information become zero, leading to poor predictor performance.
Appendix H Bounds for Jensen-Shannon Divergence (JSD) via Predictive Diversity Score (PDS)
In this section, we show that JSD is bounded quadratically below and linearly above by PDS. We first relate JSD to total variation (§ H.1), then show total variation is monotone in PDS (§ H.2), and then combine these results in § H.3.
H.1 Bounds for JSD via Total Variation (TV)
We begin by showing that JSD is bounded quadratically below and linearly above by total variation. We first introduce the setup with required definitions (§ H.1.1), then prove the proposition (§ H.1.2).
H.1.1 Setup.
Let $\{P_{\widehat{Y}_{i}\mid m}\}_{m=1}^{M}$ be distributions on $K$ classes. Define the mixture
$\bar{P}\;=\;\frac{1}{M}\sum_{m=1}^{M}P_{\widehat{Y}_{i}\mid m}.$
Definition 1 (Jensen–Shannon divergence).
$\mathrm{JSD}\!\big(\{P_{\widehat{Y}_{i}\mid m}\}\big)=\frac{1}{M}\sum_{m=1}^{M}D_{\mathrm{KL}}(P_{\widehat{Y}_{i}\mid m}\|\bar{P})=\mathcal{H}(\bar{P})-\tfrac{1}{M}\sum_{m=1}^{M}\mathcal{H}(P_{\widehat{Y}_{i}\mid m}).$
@@ H.1 Bounds for JSD via Total Variation (TV)
For distributions $P,Q$ on the same support,
$\mathrm{TV}(P,Q)=\tfrac{1}{2}\|P-Q\|_{1}.$
H.1.2 Proposition.
@@
For any $M\geq 2$ distributions $\{P_{\widehat{Y}_{i}\mid m}\}_{m=1}^{M}$ with mixture $\bar{P}$,
$\frac{2}{\ln 2}\cdot\frac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(P_{\widehat{Y}_{i}\mid m},\bar{P})^{2}\;\leq\;\mathrm{JSD}\!\big(\{P_{\widehat{Y}_{i}\mid m}\}_{m=1}^{M}\big)\;\leq\;\frac{M}{M-1}\,\log M\cdot\frac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(P_{\widehat{Y}_{i}\mid m},\bar{P}).$
@@
Lower bound. By Pinsker’s inequality (e.g. Equation 1 in (Sason, 2015)),
$D_{\mathrm{KL}}(P\|Q)\;\geq\;\frac{2}{\ln 2}\,\mathrm{TV}(P,Q)^{2}.$
Substituting $Q=\bar{P}$ and averaging over $m$ yields the lower bound.
Upper bound. Fix $m$. Write
$\bar{P}=\alpha P_{\widehat{Y}_{i}\mid m}+(1-\alpha)\zeta,\quad\alpha=\tfrac{1}{M},\quad\zeta=\tfrac{1}{M-1}\sum_{s\neq m}P_{\widehat{Y}_{i}\mid s}.$
Define $t(i)=\zeta(i)/P_{\widehat{Y}_{i}\mid m}(i)$ when $P_{\widehat{Y}_{i}\mid m}(i)>0$ (set $t(i)=+\infty$ if $P_{\widehat{Y}_{i}\mid m}(i)=0,\zeta(i)>0$). Then $\mathbb{E}_{P_{\widehat{Y}_{i}\mid m}}[t]=1$ and
$D_{\mathrm{KL}}(P_{\widehat{Y}_{i}\mid m}\|\bar{P})=\mathbb{E}_{P_{\widehat{Y}_{i}\mid m}}\!\left[-\log\big(\alpha+(1-\alpha)t\big)\right].$
Let $f(u)=-\log(\alpha+(1-\alpha)u)$, $u\geq 0$. Then $f$ is convex, decreasing, with $f(1)=0$, $f(0)=\log(1/\alpha)=\log M$. By convexity,
$f(u)\;\leq\;(1-u)f(0)=(1-u)\log M.$
@@
$D_{\mathrm{KL}}(P_{\widehat{Y}_{i}\mid m}\|\bar{P})\leq\log M\cdot\mathbb{E}_{P_{\widehat{Y}_{i}\mid m}}\!\big[(1-t)\big]\leq\log M\cdot\mathbb{E}_{P_{\widehat{Y}_{i}\mid m}}\!\big[(1-t)_{+}\big].$
@@
$\mathbb{E}_{P_{\widehat{Y}_{i}\mid m}}[(1-t)_{+}]=\sum_{i}P_{\widehat{Y}_{i}\mid m}(i)\max\{0,1-\zeta(i)/P_{\widehat{Y}_{i}\mid m}(i)\}=\sum_{i}(P_{\widehat{Y}_{i}\mid m}(i)-\zeta(i))_{+}.$
@@
$\sum_{i}(P_{\widehat{Y}_{i}\mid m}(i)-\zeta(i))_{+}=\mathrm{TV}(P_{\widehat{Y}_{i}\mid m},\zeta).$
Finally, since $\bar{P}=\alpha P_{\widehat{Y}_{i}\mid m}+(1-\alpha)\zeta$, one has
$\mathrm{TV}(P_{\widehat{Y}_{i}\mid m},\zeta)=\tfrac{M}{M-1}\,\mathrm{TV}(P_{\widehat{Y}_{i}\mid m},\bar{P}).$
@@
$D_{\mathrm{KL}}(P_{\widehat{Y}_{i}\mid m}\|\bar{P})\;\leq\;\tfrac{M}{M-1}\,\log M\cdot\mathrm{TV}(P_{\widehat{Y}_{i}\mid m},\bar{P}).$
@@
H.2 Bounds for Total Variation via Predictive Diversity Score
We next show that total variation is monotone in PDS. We introduce the setup with definitions and lemmas (§ H.2.1), then prove the proposition (§ H.2.2).
H.2.1 Setup.
Fix a class $c$. Let $X_{m}=P_{\widehat{Y}_{i}\mid m}(c)$ and $\mu=\bar{P}(c)$. Define:
@@ H.2 Bounds for Total Variation via Predictive Diversity Score
$\operatorname{PDS}\!\big(\{P_{\widehat{Y}_{i}\mid m}\}\big)=\sum^{K}_{c=1}\max_{m}P_{\widehat{Y}_{i}\mid m}(c).$
@@
Use notation from Appendix H.2.1. For each class $c$, if at most $z$ models satisfy $X_{m}>\mu$, then
@@
$E=\sum_{c=1}^{K}E_{c},\qquad U=\sum_{c=1}^{K}U_{c}=\tfrac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(P_{\widehat{Y}_{i}\mid m},\bar{P}).$
@@
H.3 Final Sandwich Inequality
Finally, we combine results from § H.1.1 and § H.2 to show that JSD is bounded quadratically below and linearly above by PDS.
@@ H.3 Final Sandwich Inequality
$\frac{2}{M^{2}\ln 2}\,(\mathrm{PDS\big(\{P_{\widehat{Y}_{i}\mid m}\}\big)}-1)^{2}\;\leq\;\mathrm{JSD}\!\big(\{P_{\widehat{Y}_{i}\mid m}\}\big)\;\leq\;\frac{M}{M-1}\log M\cdot(\mathrm{PDS\big(\{P_{\widehat{Y}_{i}\mid m}\}\big)}-1).$
@@
$U\;=\;\sum_{c=1}^{K}\frac{1}{2M}\sum_{m=1}^{M}|P_{\widehat{Y}_{i}\mid m}(c)-\bar{P}(c)|:=\frac{1}{2M}\sum_{m=1}^{M}\|P_{\widehat{Y}_{i}\mid m}-\bar{P}\|_{1}:=\frac{1}{M}\sum_{m=1}^{M}\mathrm{TV}(P_{\widehat{Y}_{i}\mid m},\bar{P}),$
where $\bar{P}=P_{\widehat{Y}_{i}}=\tfrac{1}{M}\sum_{m=1}^{M}P_{\widehat{Y}_{i}\mid m}$, and
$E\;=\;\sum_{c=1}^{K}\Big(\max_{m}P_{\widehat{Y}_{i}\mid m}(c)-\bar{P}(c)\Big)=\mathrm{PDS}-1.$
@@
Appendix I Implementation details
The results of all experiments were averaged over five runs. For the language tasks, we either used pre-computed LLM outputs downloaded from the open-llm-leaderboard (Aidar Myrzakhan, 2024) on Hugging Face or computed the outputs using lm-evaluation-harness (Gao et al., 2024). LLM evaluation was performed on a single NVIDIA H100 GPU, while parametric prediction methods (described in § 4.2.2) were trained on a single NVIDIA GTX 2080 Ti GPU. For the vision tasks, we used models from the timm library (Wightman, 2019), also evaluated on a single NVIDIA GTX 2080 Ti GPU.
Training a parametric model takes less than one minute for both vision and language domains. Details on LLM evaluation time can be seen in Table 4.
For the MMLU dataset, we observed that computing disagreement scores (as described in § 4.1.2) using all available source models led to worse DISCO performance than using only a subset of them. This can be explained by the fact that including additional, highly similar or redundant models may dilute the effective heterogeneity of the ensemble, which is crucial for DISCO as discussed in § G.0.3. Consequently, we select subsets of source models when computing disagreement scores. These subsets are obtained by randomly sampling $M$ models from the source models, where $M$ is treated as a hyperparameter and tuned jointly with other hyperparameters. For MMLU, the selected value was $M=100$. We chose random sampling here as it provides a simple and unbiased way to control ensemble size without introducing additional selection criteria.
Details for prediction methods:
  • kNN: We used $k=1$ in all experiments unless stated otherwise.
  • Random Forest (RF): Implemented using scikit-learn (Pedregosa et al., 2011) with default parameters.
  • 2-Layer MLP: Trained for 200 epochs using the AdamW optimiser (Loshchilov and Hutter, 2019) with default settings and a learning rate of 0.001. Hidden dimension: [128].
  • 3-Layer MLP: Trained for 700 epochs using the AdamW optimiser (Loshchilov and Hutter, 2019) with default settings and a learning rate of 0.001. Hidden dimensions: [128, 128].
  • Linear Regression: Implemented using scikit-learn (Pedregosa et al., 2011) with default parameters.
  • Ridge Regression: Implemented using scikit-learn (Pedregosa et al., 2011), with default parameters and regularisation weight $\lambda=10$.
  • Lasso Regression: Implemented using scikit-learn (Pedregosa et al., 2011), with default parameters and regularisation weight $\lambda=0.0001$.
  • Gradient Boosting (GB): Implemented using scikit-learn (Pedregosa et al., 2011), with default parameters and 200 base estimators.
Details for dimensionality reduction methods:
  • PCA: We used the scikit-learn implementation (Pedregosa et al., 2011) and varied the number of principal components as shown in Table 2.
  • UMAP: We used the umap-learn library (McInnes et al., 2018) and varied the number of components as specified in Table 2.
Appendix J Vision results
We introduce the setup (§J.1), describe baselines (§J.2), and present results (§J.3).
J.1 Setup